# Discount Model Search for Quality Diversity Optimization in High-Dimensional Measure Spaces

### v1

@@ 6.1. Benchmarks
Arm Repertoire (Vassiliades and Mouret, 2018) is an inverse kinematics domain where solutions are joint angle configurations of a 2D planar arm with $n=100$ joints. The measure function outputs the 2D position of the arm’s end effector. The objective indicates the variance of the $n$ joint angles.
6.2. Quality Diversity with Datasets of Measures
We propose the QDDM setting, where instead of designing measure functions, a user provides a dataset indicating their desired measure values. The defining feature of QDDM is that the user provides high-dimensional data, e.g., images, audio, or text, and the measure space $S$ is the space of such data, e.g., $S\subseteq\mathbb{R}^{k=256*256*3=196\text{\small,}608}$ if the data are $256\times 256\times 3$ RGB images.
@@ 6.2. Quality Diversity with Datasets of Measures
Triangle Arrangement (TA). We introduce two QDDM domains. First, TA builds on computational creativity domains (Tian and Ha, 2022) and involves arranging a prespecified number of triangles to create images. A solution consists of the vertices, brightness (we consider grayscale images), and alpha (transparency) of each triangle. A solution’s measure is created by rendering the triangles as a raster image. We specify desired images (measures) by sampling 1000 images from either MNIST (Lecun et al., 1998) or Fashion MNIST (Xiao et al., 2017), leading to two versions of this domain: TA (MNIST) and TA (F-MNIST). Since the images in these datasets are $28\times 28$, the measure space is 784-dimensional. Drawing from the loss function in Tian and Ha (2022), we define the distance function as Euclidean distance, so each solution is placed in the archive cell of the digit it most resembles. To make the triangle images resemble the MNIST images, we define the objective as the negative (to facilitate maximization) mean squared error between the triangle image and the archive centroid (MNIST image) to which it is assigned.
Latent Space Illumination (LSI). LSI entails exploring the latent space of a generative model to create images with diverse measures. While prior work (Fontaine et al., 2021b) considers LSI with low-dimensional measures, we consider LSI where the measures are images. As described in Sec. 1, we search for images of hikers in different landscapes. The solutions are latent vectors ($w$-space, but not $w^{+}$-space) of StyleGAN3 (Karras et al., 2021a), with images of faces output by StyleGAN3 serving as measures. Thus, the measure space is the space of $256\times 256\times 3$ images. The desired measures are specified with a dataset of 10,000 landscape images sampled from LHQ256 (Skorokhodov et al., 2021). To associate hikers with landscapes in the CVT archive, the distance function is the CLIP score (Radford et al., 2021) between the face and landscape; CLIP score is more semantically meaningful than Euclidean distance. The objective is the CLIP score between the face image and the prompt “A photo of the face of a hiker.” We also add a regularizer loss (Fontaine and Nikolaidis, 2023) to penalize latent vectors that fall outside the StyleGAN3 training distribution, and similar to TA, we reward higher alignment between faces and landscapes by adding the CLIP score between the face and the landscape to which it is assigned. We refer to this domain as LSI (Hiker).
7. Experiments
Table 1. Mean QD Score (“QD”) and Coverage (“Cov”) for each algorithm in each domain.
@@ 7. Experiments
We evaluate DMS in low- and high-dimensional measure spaces through experiments in 9 benchmarks [2D LP (Sphere), 10D LP (Sphere), 20D LP (Sphere), 50D LP (Sphere), 2D LP (Rastrigin), 10D LP (Rastrigin), 2D LP (Flat), 10D LP (Flat), Arm Repertoire] and 3 QDDM domains [TA (MNIST), TA (F-MNIST), LSI (Hiker)]. In each domain, we conduct a between-groups study with the algorithm as the independent variable: besides DMS, we consider the black-box QD algorithms described in Sec. 3: CMA-MAE, DDS, MAP-Elites (line), and MAP-Elites. We consider two dependent variables. QD Score (Pugh et al., 2016) represents overall performance by summing the objectives of all solutions in the archive, as is done in the QD objective (Sec. 2). Coverage indicates how much of the measure space has been explored by computing the percentage of archive cells that have a solution in them. Note that the objective is normalized to have a maximum value of 1 in all domains. Our hypothesis is that DMS will outperform all other algorithms in both QD Score and Coverage. We implement all algorithms with pyribs (Tjanaka et al., 2023); hyperparameters and further details are in Appendix B and E.
7.1. Analysis
Table 1 shows the results from 20 trials in the benchmark domains and 5 trials in the QDDM domains. Fig. 2 shows sample images from DMS in LSI (Hiker). Refer to Appendix A for performance plots, error bars, archive heatmaps, and discount function visualizations. We could not run DDS in the QDDM domains due to the KDE’s runtime, which grows linearly with measure space dimensionality. Visual inspection showed the results were normally distributed, but Levene’s test showed most settings were non-homoscedastic. Thus, to analyze the results, we conducted Welch’s one-way ANOVA in each domain for each dependent variable. All ANOVAs were significant ($p<0.001$; $F$ values in Appendix A), so we performed pairwise comparisons with the Games-Howell test.
@@ 7.1. Analysis
Ablation Study. We ablate the hyperparameters of DMS in Appendix C. We find that the archive learning rate $\alpha$ behaves similarly as in CMA-MAE: intermediate values enable DMS to balance optimization of the objective and exploration of the measure space, while $\alpha=0$ causes DMS to over-emphasize the objective. Meanwhile, the “empty points” are necessary for training the discount model: removing them by setting $n_{empty}=0$ causes performance to drop since the discount model takes on arbitrary values in areas of the measure space that have not been explored (Fig. 11). In contrast, setting $n_{empty}=10,100,\text{or\ }1000$ resolves this issue by “clamping down” the discount model (Fig. 8). Overall, our experimental results show how the discount model successfully guides optimization in DMS, leading to high performance across domains with different measure spaces.
@@
Computation Time. Since DMS trains a discount model, it may be more computationally expensive than other methods like CMA-MAE. To quantify this tradeoff, we recorded the wallclock time when running all the QD algorithms in our experiments. These results are summarized in Table 5 in Appendix A. We focus on comparing DMS and CMA-MAE, as their primary difference is in training the discount model. On the benchmark domains (LP and Arm Repertoire), DMS takes noticeably longer than CMA-MAE due to this training. However, on the QDDM domains (TA and LSI), this difference is much less noticeable, because the evaluation of solutions becomes the bottleneck, rather than the QD algorithm itself. Meanwhile, we note that MAP-Elites and MAP-Elites (line) are the fastest algorithms because their search operations are the least computationally expensive, since they are rooted in fixed Gaussian noise. In contrast, DMS and CMA-MAE manipulate multivariate Gaussian distributions in their CMA-ES emitters.
8. Related Work
Quality Diversity Optimization. Applications of QD include robotics (Cully et al., 2015; Huber et al., 2025; Morrison et al., 2020), drug discovery (Boige et al., 2023; Verhellen and Van den Abeele, 2020), urban planning (Galanos et al., 2021), and finance (Gašperov et al., 2025). In computer vision, QD can create diverse images by exploring the latent space of generative models, e.g., StyleGAN (Karras et al., 2021b) in our work and in (Fontaine and Nikolaidis, 2021), and Stable Diffusion (Rombach et al., 2022) in (Ding et al., 2024). Such images can form a synthetic dataset for debiasing downstream models (Chang et al., 2024). In reinforcement learning, QD can generate diverse locomotion policies (Nilsson and Cully, 2021; Pierrot et al., 2022; Batra et al., 2024; Chalumeau et al., 2023; Xue et al., 2024), and in red-teaming, QD can probe a large language model (LLM) to produce harmful outputs (Samvelyan et al., 2025).
Our work fits a growing trend of integrating models into QD. To accelerate evaluations, multiple works (Bhatt et al., 2022, 2023; Kent et al., 2025; Zhang et al., 2022; Gaier et al., 2018; Hagg et al., 2020) train surrogate models to approximate expensive objectives and/or measures. Others (Gravina et al., 2016, 2019; Salehi et al., 2021) build models that guide the creation of new solutions, especially in reinforcement learning (Tjanaka et al., 2022; Nilsson and Cully, 2021; Pierrot et al., 2022; Batra et al., 2024; Chalumeau et al., 2023; Grillotti et al., 2024). Furthermore, whereas DMS searches directly in the space of high-dimensional measures, several approaches (Cully, 2019; Paolo et al., 2020; Grillotti and Cully, 2022; McCormack and Cruz Gambardella, 2022) build models that compress high-dimensional measures into low-dimensional measures. Finally, as discussed in Sec. 6.2, we apply the manifold hypothesis (Fefferman et al., 2016) in measure space when creating the archive in QDDM domains. Prior work (Vassiliades and Mouret, 2018; Gaier et al., 2020; Rakicevic et al., 2021; Hegde et al., 2023) applies the manifold hypothesis in solution space by searching over the elite hypervolume, a low-dimensional manifold where the solutions to each QD problem are hypothesized to exist.
Related to our proposed QDDM setting, developing intuitive ways to specify measures of diversity is an active research area in QD. For example, QD through AI Feedback (QDAIF) (Bradley et al., 2024) evaluates measures by querying LLMs for feedback, while the LSI domain in prior work (Fontaine and Nikolaidis, 2021) computes the CLIP score between text prompts and generated images. QD through Human Feedback (QDHF) (Ding et al., 2024) learns measures from human preferences via a contrastive loss. Each method’s suitability depends on which user effort is easiest. For example, QDAIF excels when measures can be conveniently specified to an LLM evaluator that outputs a vector of measures. Conversely, QDAIF can be limited by the challenges of prompt engineering and the stochasticity of LLM outputs. Moreover, it may be difficult to create vector-valued measures that elicit the desired diversity, like in LSI (Hiker), where QDDM makes it easy to specify “where a hiker is located” with landscape images. A key distinction is that QDDM specifies desired measure values, while the above methods all define measure functions. If an appropriate dataset does not exist, designing a measure function may be more appropriate, since datasets require significant effort to curate. On the other hand, we believe QDDM will be applicable in many problems since datasets are abundant in machine learning.
Computational Creativity. Our TA domains draw from prior work (Tian and Ha, 2022) that arranges triangles to represent images and text prompts with evolution strategies. Similarly, various works arrange basic shapes to create artistic images (Paauw and Van den Berg, 2019; Berg et al., 2019; Cason, 2016; Fogleman, 2016). To our knowledge, QD algorithms have not been applied in this setting, but they have generated other forms of art, such as line drawings (McCormack and Cruz Gambardella, 2022) and images (Zammit et al., 2022).
Latent Space Exploration. LSI (Hiker) is an example of latent space illumination, where a QD algorithm searches for latent vectors that elicit diverse, high-performing outputs of a generative model. LSI was first introduced to generate video game levels (Fontaine et al., 2021b, a; Sarkar and Cooper, 2021; Schrum et al., 2020; Steckel and Schrum, 2021) and 2D shapes (Hagg et al., 2021). Later work (Fontaine and Nikolaidis, 2021, 2023) explored the latent space of StyleGAN (Karras et al., 2021b) to generate celebrity images with low-dimensional (2D) measures based on the CLIP score (Radford et al., 2021; Herrera-Berg, 2021). Beyond LSI, various methods aim to navigate latent spaces. Several works discover interpretable directions in GAN latent spaces, where these directions manipulate pose or facial features (Shen and Zhou, 2021; Shen et al., 2020; Voynov and Babenko, 2020; Shen et al., 2022) or transform position and scale (Plumerault et al., 2020). In single-objective optimization, prior works search the latent spaces of generative models to create video game levels (Volz et al., 2018; Merino et al., 2023; Tanabe et al., 2021) or synthetic fingerprints (Bontrager et al., 2018) that satisfy desired characteristics.
9. Conclusion
@@
Acknowledgements.
@@
Appendix A Results

### v2

@@ 6.1 Benchmarks
Arm Repertoire (Vassiliades & Mouret, 2018) is an inverse kinematics domain where solutions are joint angle configurations of a 2D planar arm with $n=100$ joints. The measure function outputs the 2D position of the arm’s end effector. The objective indicates the variance of the $n$ joint angles.
6.2 Quality Diversity with Datasets of Measures
We propose the QDDM setting, where instead of designing measure functions, a user provides a dataset indicating their desired measure values. The defining feature of QDDM is that the user provides high-dimensional data, e.g., images, audio, or text, and the measure space $S$ is the space of such data, e.g., $S\subseteq\mathbb{R}^{k=256*256*3=196\text{,}608}$ if the data are $256\times 256\times 3$ RGB images.
@@ 6.2 Quality Diversity with Datasets of Measures
Triangle Arrangement (TA). We introduce two QDDM domains. First, TA builds on computational creativity domains (Tian & Ha, 2022) and involves arranging a prespecified number of triangles to create images. A solution consists of the vertices, brightness (we consider grayscale images), and alpha (transparency) of each triangle. A solution’s measure is created by rendering the triangles as a raster image. We specify desired images (measures) by sampling 1000 images from either MNIST (Lecun et al., 1998) or Fashion MNIST (Xiao et al., 2017), leading to two versions of this domain: TA (MNIST) and TA (F-MNIST). Since the images in these datasets are $28\times 28$, the measure space is 784-dimensional. Drawing from the loss function in Tian & Ha (2022), we define the distance function as Euclidean distance, so each solution is placed in the archive cell of the digit it most resembles. To make the triangle images resemble the MNIST images, we define the objective as the negative (to facilitate maximization) mean squared error between the triangle image and the archive centroid (MNIST image) to which it is assigned.
Latent Space Illumination (LSI). LSI entails exploring the latent space of a generative model to create images with diverse measures. While prior work (Fontaine et al., 2021b) considers LSI with low-dimensional measures, we consider LSI where the measures are images. As described in Sec. 1, we search for images of hikers in different landscapes. The solutions are latent vectors ($w$-space, but not $w^{+}$-space) of StyleGAN3 (Karras et al., 2021a), with images of faces output by StyleGAN3 serving as measures. Thus, the measure space is the space of $256\times 256\times 3$ images. The desired measures are specified with a dataset of 10,000 landscape images sampled from LHQ256 (Skorokhodov et al., 2021). To associate hikers with landscapes in the CVT archive, the distance function is the CLIP score (Radford et al., 2021) between the face and landscape; CLIP score is more semantically meaningful than Euclidean distance. The objective is the CLIP score between the face image and the prompt “A photo of the face of a hiker.” We also add a regularizer loss (Fontaine & Nikolaidis, 2023) to penalize latent vectors that fall outside the StyleGAN3 training distribution, and similar to TA, we reward higher alignment between faces and landscapes by adding the CLIP score between the face and the landscape to which it is assigned. We refer to this domain as LSI (Hiker).
7 Experiments
Table 1: Mean QD Score (“QD”) and Coverage (“Cov”) for each algorithm in each domain.
@@ 7 Experiments
We evaluate DMS in low- and high-dimensional measure spaces through experiments in 9 benchmarks [2D LP (Sphere), 10D LP (Sphere), 20D LP (Sphere), 50D LP (Sphere), 2D LP (Rastrigin), 10D LP (Rastrigin), 2D LP (Flat), 10D LP (Flat), Arm Repertoire] and 3 QDDM domains [TA (MNIST), TA (F-MNIST), LSI (Hiker)]. In each domain, we conduct a between-groups study with the algorithm as the independent variable: besides DMS, we consider the black-box QD algorithms described in Sec. 3: CMA-MAE, DDS, MAP-Elites (line), and MAP-Elites. We consider two dependent variables. QD Score (Pugh et al., 2016) represents overall performance by summing the objectives of all solutions in the archive, as is done in the QD objective (Sec. 2). Coverage indicates how much of the measure space has been explored by computing the percentage of archive cells that have a solution in them. Note that the objective is normalized to have a maximum value of 1 in all domains. Our hypothesis is that DMS will outperform all other algorithms in both QD Score and Coverage. We implement all algorithms with pyribs (Tjanaka et al., 2023); hyperparameters and further details are in Appendix C and F.
7.1 Analysis
Table 1 shows the results from 20 trials in the benchmark domains and 5 trials in the QDDM domains. Fig. 2 shows sample images from DMS in LSI (Hiker). Refer to Appendix B for performance plots, error bars, archive heatmaps, and discount function visualizations. We could not run DDS in the QDDM domains due to the KDE’s runtime, which grows linearly with measure space dimensionality. Visual inspection showed the results were normally distributed, but Levene’s test showed most settings were non-homoscedastic. Thus, to analyze the results, we conducted Welch’s one-way ANOVA in each domain for each dependent variable. All ANOVAs were significant ($p<0.001$; $F$ values in Appendix B), so we performed pairwise comparisons with the Games-Howell test.
@@ 7.1 Analysis
Ablation Study. We ablate the hyperparameters of DMS in Appendix D. We find that the archive learning rate $\alpha$ behaves similarly as in CMA-MAE: intermediate values enable DMS to balance optimization of the objective and exploration of the measure space, while $\alpha=0$ causes DMS to over-emphasize the objective. Meanwhile, the “empty points” are necessary for training the discount model: removing them by setting $n_{empty}=0$ causes performance to drop since the discount model takes on arbitrary values in areas of the measure space that have not been explored (Fig. 11). In contrast, setting $n_{empty}=10,100,\text{or\ }1000$ resolves this issue by “clamping down” the discount model (Fig. 8). Overall, our experimental results show how the discount model successfully guides optimization in DMS, leading to high performance across domains with different measure spaces.
@@
Computation Time. Since DMS trains a discount model, it may be more computationally expensive than other methods like CMA-MAE. To quantify this tradeoff, we recorded the wallclock time when running all the QD algorithms in our experiments. These results are summarized in Table 2. We focus on comparing DMS and CMA-MAE, as their primary difference is in training the discount model. On the benchmark domains (LP and Arm Repertoire), DMS takes noticeably longer than CMA-MAE due to this training. However, on the QDDM domains (TA and LSI), this difference is much less noticeable, because the evaluation of solutions becomes the bottleneck, rather than the QD algorithm itself. Meanwhile, we note that MAP-Elites and MAP-Elites (line) are the fastest algorithms because their search operations are the least computationally expensive, since they are rooted in fixed Gaussian noise. In contrast, DMS and CMA-MAE manipulate multivariate Gaussian distributions in their CMA-ES emitters.
Table 2: Computation (wallclock) time (in seconds) of each algorithm in each domain. We show the mean and standard error of the mean over 20 trials for the benchmark domains and 5 trials for the QDDM domains.
2D LP (Sphere) 10D LP (Sphere) 20D LP (Sphere) 50D LP (Sphere)
DMS 397.83 ±0.37 876.33 ±44.22 2,797.61 ±6.57 7,843.98 ±228.28
CMA-MAE 121.13 ±0.10 333.92 ±1.70 1,316.87 ±22.32 4,707.92 ±86.53
DDS 1,918.68 ±22.42 3,535.78 ±1.53 1,481.18 ±30.98 2,600.24 ±78.20
MAP-Elites 39.01 ±0.21 190.25 ±2.46 873.06 ±0.35 4,523.16 ±51.12
2D LP (Rastrigin) 10D LP (Rastrigin) 2D LP (Flat) 10D LP (Flat)
DMS 707.20 ±21.31 711.08 ±30.76 683.53 ±27.73 723.34 ±32.31
CMA-MAE 182.07 ±0.36 465.80 ±3.81 165.33 ±0.25 800.12 ±0.92
DDS 1,957.29 ±0.87 3,598.86 ±2.56 1,937.41 ±1.51 3,581.79 ±2.05
MAP-Elites (line) 39.74 ±0.14 274.86 ±0.62 30.84 ±0.09 339.73 ±0.42
MAP-Elites 51.27 ±0.17 367.56 ±1.57 38.41 ±0.14 451.46 ±0.41
Arm Repertoire TA (MNIST) TA (F-MNIST) LSI (Hiker)
DMS 687.85 ±27.18 489.90 ±12.55 546.60 ±21.46 4,740.24 ±482.74
CMA-MAE 130.83 ±0.26 495.95 ±13.12 525.42 ±19.22 4,690.12 ±654.47
DDS 1,577.54 ±1.70
MAP-Elites 32.80 ±0.08 291.98 ±1.46 309.01 ±3.92 3,460.71 ±5.86
8 Conclusion
@@
9 Ethics Statement
DMS has the potential for some negative societal impacts. For example, its abilities in QDDM settings may exacerbate biases in the dataset of measures, image generator (Karras et al., 2021a), and even distance metric (Radford et al., 2021). These biases can be mitigated by carefully managing all datasets, including that used to train the generator (Perera & Patel, 2023; Maluleke et al., 2022). In general, QD can also debias models by generating balanced datasets (Chang et al., 2024). In addition, using QDDM methods to drive generative search towards specific artistic styles defined in the measure space can discount unique styles created by human artists (Shan et al., 2023). Alternative objective functions (Tseng et al., 2021) may enable finding solutions that reflect the dataset of measures without mimicking artists’ creativity.
10 Reproducibility Statement
We have included details of DMS in Sec. 5, including pseudocode in Algorithm 1. We include details of our domains and experimental setup in Sec. 6 and Sec. 7, along with hyperparameters in Appendix C, extended domain details in Appendix E, and further implementation details in Appendix F. We have released our code at https://github.com/icaros-usc/discount-models. Inside the code repository, a README file lists instructions for setting up the environment and reproducing all experiments in this paper.
Acknowledgments
@@
Appendix A Related Work
Quality Diversity Optimization. Applications of QD include robotics (Cully et al., 2015; Huber et al., 2025; Morrison et al., 2020), drug discovery (Boige et al., 2023; Verhellen & Van den Abeele, 2020), urban planning (Galanos et al., 2021), and finance (Gašperov et al., 2025). In computer vision, QD can create diverse images by exploring the latent space of generative models, e.g., StyleGAN (Karras et al., 2021b) in our work and in (Fontaine & Nikolaidis, 2021), and Stable Diffusion (Rombach et al., 2022) in (Ding et al., 2024). Such images can form a synthetic dataset for debiasing downstream models (Chang et al., 2024). In reinforcement learning, QD can generate diverse locomotion policies (Nilsson & Cully, 2021; Pierrot et al., 2022; Batra et al., 2024; Chalumeau et al., 2023; Xue et al., 2024), and in red-teaming, QD can probe a large language model (LLM) to produce harmful outputs (Samvelyan et al., 2025).
Our work fits a growing trend of integrating models into QD. To accelerate evaluations, multiple works (Bhatt et al., 2022; 2023; Kent et al., 2025; Zhang et al., 2022; Gaier et al., 2018; Hagg et al., 2020) train surrogate models to approximate expensive objectives and/or measures. Others (Gravina et al., 2016; 2019; Salehi et al., 2021) build models that guide the creation of new solutions, especially in reinforcement learning (Tjanaka et al., 2022; Nilsson & Cully, 2021; Pierrot et al., 2022; Batra et al., 2024; Chalumeau et al., 2023; Grillotti et al., 2024). Furthermore, whereas DMS searches directly in the space of high-dimensional measures, several approaches (Cully, 2019; Paolo et al., 2020; Grillotti & Cully, 2022; McCormack & Cruz Gambardella, 2022) build models that compress high-dimensional measures into low-dimensional measures. Finally, as discussed in Sec. 6.2, we apply the manifold hypothesis (Fefferman et al., 2016) in measure space when creating the archive in QDDM domains. Prior work (Vassiliades & Mouret, 2018; Gaier et al., 2020; Rakicevic et al., 2021; Hegde et al., 2023) applies the manifold hypothesis in solution space by searching over the elite hypervolume, a low-dimensional manifold where the solutions to each QD problem are hypothesized to exist.
Related to our proposed QDDM setting, developing intuitive ways to specify measures of diversity is an active research area in QD. For example, QD through AI Feedback (QDAIF) (Bradley et al., 2024) evaluates measures by querying LLMs for feedback, while the LSI domain in prior work (Fontaine & Nikolaidis, 2021) computes the CLIP score between text prompts and generated images. QD through Human Feedback (QDHF) (Ding et al., 2024) learns measures from human preferences via a contrastive loss. Each method’s suitability depends on which user effort is easiest. For example, QDAIF excels when measures can be conveniently specified to an LLM evaluator that outputs a vector of measures. Conversely, QDAIF can be limited by the challenges of prompt engineering and the stochasticity of LLM outputs. Moreover, it may be difficult to create vector-valued measures that elicit the desired diversity, like in LSI (Hiker), where QDDM makes it easy to specify “where a hiker is located” with landscape images. A key distinction is that QDDM specifies desired measure values, while the above methods all define measure functions. If an appropriate dataset does not exist, designing a measure function may be more appropriate, since datasets require significant effort to curate. On the other hand, we believe QDDM will be applicable in many problems since datasets are abundant in machine learning.
Computational Creativity. Our TA domains draw from prior work (Tian & Ha, 2022) that arranges triangles to represent images and text prompts with evolution strategies. Similarly, various works arrange basic shapes to create artistic images (Paauw & Van den Berg, 2019; Berg et al., 2019; Cason, 2016; Fogleman, 2016). To our knowledge, QD algorithms have not been applied in this setting, but they have generated other forms of art, such as line drawings (McCormack & Cruz Gambardella, 2022) and images (Zammit et al., 2022).
Latent Space Exploration. LSI (Hiker) is an example of latent space illumination, where a QD algorithm searches for latent vectors that elicit diverse, high-performing outputs of a generative model. LSI was first introduced to generate video game levels (Fontaine et al., 2021b; a; Sarkar & Cooper, 2021; Schrum et al., 2020; Steckel & Schrum, 2021) and 2D shapes (Hagg et al., 2021). Later work (Fontaine & Nikolaidis, 2021; 2023) explored the latent space of StyleGAN (Karras et al., 2021b) to generate celebrity images with low-dimensional (2D) measures based on the CLIP score (Radford et al., 2021; Herrera-Berg, 2021). Beyond LSI, various methods aim to navigate latent spaces. Several works discover interpretable directions in GAN latent spaces, where these directions manipulate pose or facial features (Shen & Zhou, 2021; Shen et al., 2020; Voynov & Babenko, 2020; Shen et al., 2022) or transform position and scale (Plumerault et al., 2020). In single-objective optimization, prior works search the latent spaces of generative models to create video game levels (Volz et al., 2018; Merino et al., 2023; Tanabe et al., 2021) or synthetic fingerprints (Bontrager et al., 2018) that satisfy desired characteristics.
Appendix B Results